### v1

# An “opinion reproduction number” for infodemics in a bounded-confidence content-spreading process on networks

Abstract
We study the spreading dynamics of content on networks. To do this, we use a model in which content spreads through a bounded-confidence mechanism. In a bounded-confidence model (BCM) of opinion dynamics, the agents of a network have continuous-valued opinions, which they adjust when they interact with agents whose opinions are sufficiently close to theirs. The employed content-spread model introduces a twist into BCMs by using bounded confidence for the content spread itself. To study the spread of content, we define an analogue of the basic reproduction number from disease dynamics that we call an opinion reproduction number. A critical value of the opinion reproduction number indicates whether or not there is an “infodemic” (i.e., a large content-spreading cascade) of content that reflects a particular opinion. By determining this critical value, one can determine whether or not an opinion will die off or propagate widely as a cascade in a population of agents. Using configuration-model networks, we quantify the size and shape of content dissemination using a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics. We find that content spreads most widely when the agents have large expected mean degree or large receptiveness to content. When the amount of content spread only slightly exceeds the critical opinion reproduction number (i.e., the infodemic threshold), there can be longer dissemination trees than when the expected mean degree or receptiveness is larger, even though the total number of content shares is smaller.
Although most content does not spread far on social media, occasionally some piece of content “goes viral” and rapidly reaches many people Goel et al. (2016); Juul and Ugander (2021). When content with misinformation or disinformation spreads widely, it has become common to expand on analogies with disease spread by stating that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021), which is a portmanteau of the words “information” and “epidemic”. In the present paper, we examine infodemics in a model of content spread that uses a bounded-confidence mechanism. We extend the analogy between the spread of content and the spread of infectious diseases by defining an opinion reproduction number, which is inspired by the basic reproduction number of disease dynamics Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). By examining whether or not the amount of opinion spread is below or above a critical opinion reproduction number (i.e., an infodemic threshold), one can examine whether content dies out or goes viral. This approach complements branching-process approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021). In our bounded-confidence model of content spread, we quantify the size and shape of content dissemination using a variety of summary statistics, and we illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks.
I Introduction
Given the enormous scale and impact of human interactions on social media, it is critical to study the collective dynamics that arise in these systems Bak-Coleman et al. (2021). Content on social media can take many forms, including scientific facts, other forms of information, hyperlinks to Web pages, pictures, memes, and misinformation and disinformation Guille et al. (2013); Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022). Once created, content can then be magnified by human users, bots, and other accounts, creating an online ecosystem that is filled with much misinformation and disinformation Starbird (2019); Yang et al. (2023), even with intentions to share only accurate information Pennycook et al. (2021). Content sharing by people is ubiquitous and important for social connection, and such sharing can help create and reinforce shared understanding Baek et al. (2023). Most content does not spread to many people, but a small proportion of online content spreads very far (i.e., it can “go viral”) in large content-spreading cascades Goel et al. (2016); Juul and Ugander (2021). A variety of other social phenomena, including emotions and behaviors, can spread analogously as “social contagions” on networks Aral, Muchnik, and Sundararajan (2009); Christakis and Fowler (2013); Borge-Holthoefer et al. (2013); Yu et al. (2020); Pröllochs, Bär, and Feuerriegel (2021).
When a piece of misinformation or disinformation spreads very widely, it has become reasonably common to state that there is an infodemic Eysenbach (2002); Briand et al. (2021); Zielinski (2021). The World Health Organization (WHO) gives a much more specific definition of the term “infodemic” (a portmanteau of the words “information” and “epidemic”) in the context of infectious diseases World Health Organization (2023); Calleja et al. (2021): “An infodemic is too much information including false or misleading information in digital and physical environments during a disease outbreak.” Using both the rigid WHO definition and the looser notion of particularly low-quality content going viral, the COVID-19 pandemic has also been an infodemic Zarocostas (2020); Cinelli et al. (2020); Gallotti et al. (2020). Analogously to the spread of infectious diseases “superspreader” social-media accounts have played an important role in the COVID-19 infodemic Yang et al. (2021).
Developing a thorough understanding of misinformation, disinformation, and their impact requires a broad view of the problem of “fake news” that also encompasses proper understanding of misinformation and its effects Lazer et al. (2018); Watts, Rothschild, and Mobius (2021). This view needs to encompass disinformation (which is intentionally incorrect), misinformation (which is incorrect, but perhaps unintentionally), biased and misleading information (which may not be factually incorrect), and the production and amplification of such content, including by mainstream news organizations. There has been much empirical research on misinformation in areas such as computational social science and allied disciplines Muhammed T and Mathew (2022). Modeling efforts can play an important role in mitigating the harmful effects of misinformation and disinformation Rabb et al. (2022). In particular, Juul and Ugander Juul and Ugander (2021) suggested that focusing on reducing the “infectiousness” of information in concert with theoretical analysis of spreading processes may be very helpful for limiting the spread of misinformation and disinformation.
One strategy to gain insight into the mechanisms that underlie observations in social-media systems is by studying mathematical models of opinion dynamics on networks Noorazar et al. (2020); Noorazar (2020); Peralta, Kertész, and Iñiguez (2022). Opinion models take a variety of forms. The nodes of a network represent agents, and the edges between agents indicate social and/or communication ties between agents. The opinions of the agents can take either discrete values (e.g., $+1$ or $-1$) or continuous values (e.g., in the interval $[-1,1]$, with $-1$ the most liberal opinion and $+1$ the most conservative opinion). When two (or more) adjacent agents interact, one or more of them updates their opinion according to some rule. One popular type of model of opinion dynamics are bounded-confidence models (BCMs) Lorenz (2007); Meng, Van Gorder, and Porter (2018); Bernardo et al. (2024), in which agents have continuous-valued opinions and interacting agents compromise their opinions by some amount if and only if their opinions are sufficiently close to each other. There have been numerous studies of BCMs, which have been generalized in many ways. Recent studies have incorporated phenomena such as media outlets with fixed opinions Brooks and Porter (2020), polyadic interactions (i.e., interactions between three or more agents) Hickok et al. (2022); Wang (2022), noise Goddard et al. (2022), asymmetric confidence intervals Bernardo, Altafini, and Vasca (2022), cost of opinion change Schawe and Hernández (2020), agents with heterogeneous activity levels Li and Porter (2023), smooth interaction kernels (in the form of sigmoidal functions) to describe how agents influence each other Brooks, Chodrow, and Porter (2023), opinion repulsion Kann and Feng (2023), homophilic adaptivity of network structure Kan, Feng, and Porter (2023), preserving connectivity of the interaction network Li, Motsch, and Weber (2020), and adaptive confidence bounds Li, Luo, and Porter (2023).
Another line of modeling research is the study of spreading dynamics on networks Borge-Holthoefer et al. (2013); Porter and Gleeson (2016). Research on spreading dynamics often uses ideas from percolation theory Sahimi (2023). The study of spreading processes on networks includes research on social dynamics Lehmann and Ahn (2018), disease dynamics Pastor-Satorras et al. (2015), and how they affect each other Bedson et al. (2021). Much work on so-called “social contagions” has been inspired by the many decades of work on biological contagions, although researchers argue about whether and how much social phenomena spread in a manner that resembles disease spread Weng, Menczer, and Ahn (2013); Lerman (2016); Hui et al. (2018); Olsson and Galesic (2023); St-Onge, Hébert-Dufresne, and Allard (2024). However, despite the differences between disease spread and the spread of information and other content, it it worthwhile to explore parallels between them. Whether one is considering a disease or a piece of online content, some things spread very far (i.e., they go viral) before dissipating and others die out very quickly Juul and Ugander (2021). In some models of social phenomena, such as simplistic cascade models and anything else that takes the form of a branching process Noël et al. (2012), one can study whether the spread of content tends to magnify or contract locally as a function of time by computing branching numbers Gleeson et al. (2021), which indicate the mean number of offspring (e.g., the mean number of reposts as a function of time) of a piece of content. This is a simplification of empirical content spread on social media, where it is common to trace the spreading paths of posts, post boosting, and post commenting and quoting using dissemination trees Gomez Rodriguez, Leskovec, and Krause (2010); Guille et al. (2013); Goel et al. (2016); Kozitsin (2023). Analogously to studies of content spreading in social systems, when researchers study compartmental models of disease spread, it is traditional to calculate reproduction numbers to examine whether or not a disease will die out Brauer, Castillo-Chavez, and Feng (2019). The most standard type of reproduction number is the basic reproduction number, which measures how many infections occur, on average, from one infected node in a population in which all other nodes are susceptible to infection Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019). Compartmental models have also been employed in the study of opinion dynamics, and one can then calculate a basic reproduction number to examine whether or not content goes viral Xu et al. (2022). In the study of disease spread, branching-process theory has been used in concert with maximum-likelihood estimation to estimate basic reproduction numbers from data Blumberg and Lloyd-Smith (2013).
In the present paper, we combine ideas from opinion models and percolation-inspired spreading models to study the mean “infectiousness” of content in a content-spread model that uses a bounded-confidence mechanism for the spreading. To analyze the structure of content spreading in our model, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics, and we use this opinion reproduction number and its associated critical value (which we call the “infodemic threshold”) to examine when an infodemic occurs in a network. We use generating functions to derive an analytical expression for the opinion reproduction number on configuration-model networks in the limit of infinitely many nodes. We then compare this analytical result with numerical simulations. We quantify the spread of content by calculating the total number of pieces of content, the longest adoption paths, the widths of dissemination trees, and structural virality.
Our paper proceeds as follows. In section II, we give a brief introduction to bounded-confidence models of opinion dynamics. In section III, we describe a content-spread process that draws some inspiration from percolation theory and uses a bounded-confidence mechanism for spreading. In section IV, we analyze the conditions in which there is a large content-spreading cascade (i.e., an “infodemic”). To allow us to forecast whether or not an infodemic occurs in a network, we define an opinion-dynamics analogue of the basic reproduction number from disease dynamics. In section V, we quantify the total number of shared pieces of content on a finite-size network. In section VI, we compute a variety of summary statistics on dissemination trees to illustrate how network structure and spreading-model parameters affect these statistics in configuration-model networks. In section VII, we conclude and discuss our results. In section VIII, Mason writes a few words about David Campbell and wishes him a wonderful 80th birthday. The code to reproduce our numerical simulations is available at https://github.com/hzinnbrooks/bounded-confidence-spreading-process.
II Bounded-confidence models on networks
Bounded-confidence models (BCMs) Lorenz (2007); Proskurnikov and Tempo (2018); Noorazar et al. (2020); Noorazar (2020) of opinion dynamics include both consensus-seeking behavior and preferences for similar views (through “selective exposure”). BCMs have been used by many researchers to study social influence and group dynamics (including consensus, polarization, fragmentation, and other phenomena) Bernardo et al. (2024).
We model a social network as a graph $G(V,\mathcal{E})$, where $V$ (with $|V|=N$, so $N$ is the “size” of the network) is the set of vertices (i.e., nodes) and $\mathcal{E}\subseteq V\times V$ is the set of edges. Each node represents an agent in some population, and each edge encodes a social and/or communication tie between two agents. Node $i$ is adjacent to node $j$ if there is an edge between $i$ and $j$. We assume that $G$ is simple, so it has no self-edges or multi-edges. The graph $G$ has an associated $N\times N$ adjacency matrix $A$, where $A_{ij}=1$ if $i$ is adjacent to $j$ (and vice versa) and $A_{ij}=0$ otherwise. One can consider weighted edges to model the strengths of ties between nodes. In this case, there is also a weight matrix $W$ with entries $w_{ij}\in\mathbb{R}_{\geq 0}$. If desired, one can also consider directed networks and other generalizations of graphs. See Wasserman and Faust (1994); Newman (2018); Bullo (2022); Brooks (2023) for books and reviews about networks.
Suppose that each node $i$ has a time-dependent opinion (i.e., ideology) $x_{i}(t)$, which takes continuous values on some domain that we call the opinion space. For example, the opinion space can be the real line $\mathbb{R}$, a closed and bounded interval $[a,b]\subset\mathbb{R}$, or a subset of a higher-dimensional space $\mathbb{R}^{k}$. In a BCM, the key feature is that agents are receptive only to other agents that are within a distance $c$ of their current opinion. This parameter $c$ is called the confidence bound. If $d(x_{i}(t),x_{j}(t))<c$, where $d$ is a metric on the opinion space, we say that agent $i$ is receptive to agent $j$ (and vice versa). Agents that are receptive to each other influence each other’s opinions through an update rule. It is also possible to consider BCMs with asymmetric receptiveness, but we examine only symmetric situations.
III A content-spreading process with a bounded-confidence mechanism
We study a content-spreading model with a bounded-confidence mechanism for adopting an opinion and sharing content. However, in our model, receptive agents simply adopt the opinion that is espoused by content, rather than compromising their opinion as in a standard BCM.
The bounded-confidence models in section II describes the time-dependent opinions of the agents of a network. However, in real life, one does not usually have easy access to numerical values of peoples opinions. One instead typically observes the content that is spread on social networks. Naturally, such content often reflects an underlying opinion or ideological belief. There is much empirical work on quantifying the dissemination and virality of content on social-media platforms like $\mathbb{X}$ (which is the platform formerly known as Twitter) Kümpel, Karnowski, and Keyling (2015); Muhammed T and Mathew (2022), and branching processes provide a natural model of content spread on social media. See, for example, Lerman and Ghosh (2010); Guille et al. (2013); Gleeson et al. (2014); Vosoughi, Roy, and Aral (2018); Gleeson et al. (2021).
We seek to track the spread of fixed content with a fixed ideology (i.e., opinion) $x_{0}$. In other words, the spreading process has no mutation of ideas or editorializing. In the context of social media, one can interpret the fixed-content assumption as accounts only reposting or sharing content, without any additions or changes. By contrast, real social-media systems have a rich ecosystem of content evolution Adamic et al. (2016).
As described in section II, we model a social network as a simple, undirected graph $G(V,\mathcal{E})$ with $|V|=N$ nodes. Let $A$ denote the graph’s associated $N\times N$ adjacency matrix. The adjacency-matrix entry $A_{ij}=1$ if there is an edge between nodes $i$ and $j$; otherwise, $A_{ij}=1$. The spread of content with opinion $x_{0}$ begins at a source node, which we select uniformly at random from $V$ and label as node 00. We initialize all other nodes $i\in\{1,\ldots,N-1\}$ in $V$ with a state $x_{i}$ that we draw from a distribution $\phi(x)$. We require $\phi(x)$ to be Riemann integrable.
Given an opinion space $\Omega$, we refer to the opinion $x_{0}\in\Omega$ of the content as the content state. Additionally, we refer to the state $x_{i}\in\Omega$ of node $i\in\{1,\ldots,N-1\}$ as its opinion. In the present paper, we focus on $\Omega=[0,1]$, although in Figure 10 we show a situation with $\Omega=\mathbb{R}$. Let $c\in[0,1]$ denote the receptiveness parameter, which is akin to the confidence bound of a typical BCM. In practice, in the present paper, we consider $c\in(0,1/2]$, so that agents are not receptive to agents with overly different opinions. The spread of content with opinion $x_{0}$ begins at node 00. The content spreads through a percolation-inspired process Sahimi (2023) via an update rule that is based on a bounded-confidence mechanism (see algorithm 1). In fig. 1, we show a schematic of our algorithm on a small network.
Algorithm 1 Our algorithm for the spread of content on a network.
Input: model parameters $c,\ x_{0},\ \phi(x),\ N$, and the distribution of the graph degree sequence
Output: Dissemination tree $G_{0}$
[…]
     for $i$ in active nodes do
         neighbors =$=\{j\ |\ A_{ij}=1\}$
         for $j$ in neighbors do
$x_{j}=\begin{cases}x_{i}\text{ (and $i$ is appended to {next nodes})}\,,&\text{if }|x_{j}-x_{0}|<c\\ x_{j}\,,&\text{otherwise}\end{cases}$
         end for
         for $k$ in next nodes do
              $A_{ik}=0$ for all $i$
         end for
     end for
     active nodes $=$ next nodes
     Add next nodes to $G_{0}$
[…]
end while
Figure 1: A schematic illustration of our content-spreading process (see Algorithm 1). In the top row, we show consecutive update steps of Algorithm 1. We initialize a graph at time $t=0$ with one active node (in pink), which we label as node 00. We suppose that the root node 00 has content state $x_{0}=0.1$. We initialize all other nodes $j$ to have opinions $x_{j}\in(0,1)$. In this example, the receptiveness parameter is $c=0.35$. At each time step, any neighbors $j$ of the previously active node that satisfy $|x_{j}-x_{0}|<c$ become active (i.e., turn pink). We thus add them to the dissemination tree $G_{0}$. In this example, Algorithm 1 terminates after 4 steps. In the bottom row, we show the resulting dissemination tree $G_{0}$ in orange, with the root node outlined.
To study content spread on a network, we construct a dissemination tree $G_{0}$ with root node 00 to track the spreading process Goel et al. (2016); Oh and Porter (2018). Content spreads as follows. Suppose that content state $x_{0}$ starts at node 00. We look at the set $\mathcal{N}_{i}$ of neighbors $j$ of node $i$ and select all $j\in\mathcal{N}_{i}$ that satisfy $|x_{j}-x_{0}|<c$. These neighbors of $i$ spread the content (i.e., they become “activated” with the content and change their opinion $x_{j}$ to be equal to the content state $x_{0}$) if the distance between their opinion $x_{j}$ and the content state $x_{0}$ is sufficiently small. We represent this activation by adding these nodes to the dissemination tree $G_{0}$ as child nodes of node $i$. We then continue this process for each child node of the current dissemination tree until there are no available neighbors that are left to activate. In other words, we continue this process until either all nodes are activated or until there are no remaining neighbors of an active node.
Using Algorithm 1, it is guaranteed that $G_{0}$ is a tree because a node activates the first time that it encounters the content of one of its neighbors and can only be activated once. When two activated nodes share a receptive neighbor, we choose the activated node with the smaller index as the parent node for the newly activated neighbor. Modifications of this algorithm can result in a directed acyclic graph instead of a tree. It is worthwhile to examine such modifications in future work (especially if considering the effects of competing social contagions Min and San Miguel (2018)).
IV Forecasting an “infodemic” on a large network
We now examine content spread on a network with infinitely many nodes. This situation approximates content spread on networks with many nodes (i.e., for large $N$), particularly in the early stages of content spread.
We follow the analysis in Noël et al. (2009), who examined a model of infectious-disease spread. We use generating functions to write down the expected number of neighbors that spread the source node’s content. We refer to this expected number of neighbors as the opinion reproduction number $R$. The opinion reproduction number, whose critical value $R=1$ is its associated infodemic threshold, is akin to the basic reproduction number $R_{0}$ in models of disease spread Kiss, Miller, and Simon (2017); Brauer, Castillo-Chavez, and Feng (2019).
IV.1 Analysis
Suppose that content spreads on a configuration-model network Fosdick et al. (2018) with degree distribution $p_{k}$. We generate each configuration-model network using stub matching, without removing self-edges or multi-edges. Let $q_{k}$ denote its associated excess degree distribution. The generating function $g_{0}$ of the degree distribution and the generating function $g_{1}$ of the excess degree distribution $q_{k}$ are
$g_{0}(z)=\sum_{k=0}^{\infty}p_{k}z^{k}\,,$ (1)
$g_{1}(z)=\sum_{k=0}^{\infty}q_{k}z^{k}\,.$ (2)
Given a node with $k$ neighbors, we want to write down the probability that $l$ of these $k$ neighbors spread a piece of content. To be able to do this, we need to calculate the probability that content is spread along one edge. This single-edge transmission probability depends on both the content state $x_{0}$, the receptiveness parameter $c$, and the distribution of the initial opinions $x_{i}$ (with $i\in\{1,\ldots,N\}$).
Suppose that we draw the initial opinions $x_{i}$ from a distribution with probability density function $\phi(x)$. The probability of content spread (i.e., transmission) along an edge depends on the probability of drawing a value $x_{i}$ that lies within distance $c$ of $x_{0}$. The single-edge transmission probability is
[…]
One common choice of $\phi(x)$ is to draw the initial node opinions from the uniform distribution on the interval $(0,1)$. The probability of drawing an opinion $x_{i}$ within distance $c$ of $x_{0}$ is then $2c$, unless $x_{0}$ is within $c$ of the boundaries of $(0,1)$. This yields
$s(x_{0},c)=\begin{cases}c+x_{0}\,,&x_{0}\leq c\\ 2c\,,&x_{0}\in(c,1-c]\\ 1+c-x_{0}\,,&x_{0}>1-c\,.\end{cases}$ (4)
Another common choice of $\phi(x)$ is to draw the initial node opinions from a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$ (where the opinion space is $\mathbb{R}$). This yields
$s(x_{0},c)=\frac{1}{2}\left[\mathrm{erf}\left(\frac{x_{0}+c-\mu}{\sigma\sqrt{2}}\right)-\mathrm{erf}\left(\frac{x_{0}-c-\mu}{\sigma\sqrt{2}}\right)\right]\,,$ (5)
[…]
Using the single-edge transmission probability [see eq. 3], we see that the probability that $l$ of $k$ neighbors spread the content is
$p(l|k)=\binom{k}{l}\left(s(x_{0},c)\right)^{l}\left(1-s(x_{0},c)\right)^{k-l}\,.$ (6)
Therefore, the probability generating function of the distribution of the number of content shares that originate from the source node 00 is
$\sum_{l=0}^{\infty}\sum_{k=l}^{\infty}p_{k}p(l|k)z^{l}$ $=\sum_{l=0}^{\infty}\sum_{k=l}^{\infty}p_{k}\binom{k}{l}\left(s(x_{0},c)\right)^{l}\left(1-s(x_{0},c)\right)^{k-l}z^{l}$
$=\sum_{k=0}^{\infty}p_{k}\sum_{l=0}^{k}\binom{k}{l}\left(s(x_{0},c)\right)^{l}\left(1-s(x_{0},c)\right)^{k-l}z^{l}$
$=\sum_{k=0}^{\infty}p_{k}\left(1+(z-1)s(x_{0},c)\right)^{k}$
$=g_{0}\left(1+(z-1)s(x_{0},c)\right)\,.$ (7)
To write down the probability generating function for the number of second neighbors (i.e., the neighbors of neighbors) of a source node in a dissemination tree, we use a similar argument, but now we use the excess degree distribution instead of the degree distribution. We obtain
$g_{1}\left(1+(z-1)s(x_{0},c)\right)\,.$ (8)
If a network $G$ is infinite (or at least large enough so that we can neglect finite-size effects), nodes cannot be activated more than once, and the content spread yields a directed acyclic graph (i.e., the dissemination tree), it follows that $g_{a}\left(1+(z-1)s(x_{0},c)\right)=g_{1}\left(1+(z-1)s(x_{0},c)\right)$ is the probability generating function for the number of nodes that spread a piece of content that they received directly from a single node at level $a$ of the dissemination tree. Importantly, we are relying on the assumption that the number of times that a piece of content is spread is much smaller than the total number of nodes of a network. We expect our approximation to be reasonable in the first few generations of a dissemination tree, which describe the early stages of content spread.
In the onset of content dissemination, the expected number of “grandchildren” (i.e., second neighbors) in a dissemination tree that starts from a node that spreads the content at level $a$ is the opinion reproduction number
$R=\frac{\mathrm{d}}{\mathrm{d}z}g_{a}\left(1+(z-1)s(x_{0},c)\right)\bigg{|}_{z=1}$ $=s(x_{0},c)g_{1}^{\prime}(1)$
$=s(x_{0},c)\frac{g_{0}^{\prime\prime}(1)}{g_{0}^{\prime}(1)}\,,$ (9)
where the prime ${}^{\prime}$ denotes differentiation with respect to $z$. When the number of content shares increases from one level to the next, the opinion reproduction number $R$ (which is analogous to the basic reproduction number $R_{0}$ in models of infectious-disease spread Kiss, Miller, and Simon (2017)) is larger than $1$. Similarly, a decrease in number of content shares from one level to the next corresponds to $R<1$. The critical value $R=1$ is the infodemic threshold. This calculation is reminiscent of branching-processes approaches that examine whether or not content spread tends to locally magnify or contract with time Gleeson et al. (2021).
We now calculate $R$ explicitly for an example in which we draw the initial node opinions uniformly at random from the interval $(0,1)$. In a configuration-model network with a Poisson degree distribution with mean $\lambda$, the opinion reproduction number is
$R=s(x_{0},c)\frac{\lambda^{2}}{\lambda}=s(x_{0},c)\lambda$
because $g_{0}(z)=e^{-\lambda(1-z)}$. The transition value that determines whether content spread increases locally or decreases locally is
$R=\begin{cases}\frac{1}{\lambda}+x_{0}\,,&x_{0}\leq c\\ \frac{1}{2\lambda}\,,&c<x_{0}<1-c\\ \frac{1}{\lambda}-1+x_{0}\,,&x_{0}\geq 1-c\,.\end{cases}$
When the receptiveness parameter $c>R$, the expected number of content shares increases from one level to the next in the early stages of content spread, so we expect the content to take hold in the population. Unsurprisingly, the condition $c>R$ is satisfied most easily for larger receptive parameter $c$ and larger mean degree $\lambda$. The content state $x_{0}$ has less impact than $c$ and the node degree, except when the content state is near 00 or $1$.
IV.2 Simulations
We now compare our analytical results from section IV.1 to numerical simulations of our content-spread model. We simulate the model on 10000-node configuration-model networks with degree sequences that we determine using a Poisson distribution with mean $\lambda=5$. We draw the initial opinions uniformly at random from the interval $(0,1)$. In fig. 2, we show a scenario in which our analysis indicates that we are above the infodemic threshold (i.e, $R>1$). In fig. 3, we show a scenario in which our analysis indicates that we are below the infodemic threshold (i.e., $R<1$). In this computation, the mean total number of content shares is very small. (It is about 10.)
In fig. 4, we show a phase diagram that summarizes when our content-spreading model experiences experiences an infodemic for different values of the content state $x_{0}$ and receptiveness parameter $c$. We show the infodemic threshold, which we obtain using the condition $R=1$, from our analysis in orange.
Figure 2: A comparison of numerical simulations of our content-spreading model with our analytical predictions in an infodemic situation (i.e., when $R>1$). For each numerical simulation, we generate a 10000-node configuration-model network with a degree sequence that we determine using a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.2$. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. The solid purple curve gives the total number of content shares up to and including time $t$ averaged over $100$ realizations. In each realization, we draw a new degree sequence and a new set of node opinions. The light purple shaded region indicates the standard deviation for these $100$ realizations. The dashed orange curve gives our analytical approximation.
Figure 3: A comparison of numerical simulations of our content-spreading model with our analytical predictions in a situation without an infodemic (i.e., when $R<1$). For each numerical simulation, we generate a 10000-node configuration-model network with a degree sequence that we determine using a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.05$. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. The solid purple curve gives the total number of content shares up to and including time $t$ averaged over $100$ realizations. In each realization, we draw a new degree sequence and a new set of node opinions. The light purple shaded region indicates the standard deviation for these $100$ realizations. The dashed orange curve gives our analytical approximation.
Figure 4: The $(x_{0},c)$ phase diagram for 10000-node configuration-model networks with degree sequences that we generate using a Poisson distribution with mean $\lambda=5$. We vary both the content state ($x_{0}$, horizontal axis) and the receptiveness parameter ($c$, vertical axis). The shading of each square indicates the mean, across 100 trials, of the total number of content shares as a proportion of the total number of nodes, with lighter shades indicating that more nodes share the content. In each realization, we choose $5$ source nodes uniformly at random to seed with content state $x_{0}$. In each realization, we draw a new degree sequence and a new set of node opinions. The orange solid curve is the predicted opinion reproduction threshold $R$ from our analysis.
V Quantifying content spread
Now that we have predicted the onset of infodemics in our bounded-confidence content-spreading model on configuration-model networks, we measure the sizes (i.e., the numbers of nodes) in the resulting dissemination trees. The size of a dissemination tree equals the total number of content shares. Studying dissemination trees allows us to explore the long-term behavior of our content-spreading model. These calculations thereby complement our examination of early-stage spreading in section IV.
V.1 Analysis of the total number of content shares
Given the number of nodes and the degree distribution of a configuration-model graph $G$, the content state $x_{0}$, the receptiveness parameter $c$, and the distribution $\phi(x)$ of opinions, we again use generating functions (as in section IV.1) to estimate the total number of content shares (i.e., the size) of a dissemination tree $G_{0}$. This analysis treats content spread as a percolation process. See Newman (2018) for a detailed description of this approach.
Let $S$ be the fraction of a network’s nodes that spread a piece of content. There are thus a total of $SN$ content shares. Node opinions are independent of connectivity in our content-spreading model, so the probability that a node is in the “spreading set” $\mathcal{S}$ equals the probability that it is receptive to the content multiplied by the probability that it belongs to the connected component with the source node. Let $u$ be the mean probability of not being connected to this component via a particular neighbor. We then have
[…]
$=s(x_{0},c)\left(1-g_{0}(u)\right)\,,$ (10)
where we recall that $p_{k}$ is the probability a node has degree $k$ and $g_{0}(u)$ is the generating function of the degree distribution of the network.
To calculate $S$, we need to determine $u$. A particular node $i$ not being connected to the spreading set $\mathcal{S}$ via a particular node $j$ can occur either because it is not receptive to the content or because it is not in the same component as the content. In other words, it has been isolated from the content spread because of other non-receptive nodes. For a node $j$ with $k$ edges aside from the one to node $i$, this occurs with probability $1-s(x_{0},c)+s(x_{0},c)u^{k}$, Therefore,
$u$ $=\sum_{k=0}^{\infty}q_{k}\left(1-s(x_{0},c)+s(x_{0},c)u^{k}\right)$
$=1-s(x_{0},c)+s(x_{0},c)g_{1}(u)\,,$ (11)
where $q_{k}$ again denotes the probability of having $k$ neighbors other than the spreading edge and $g_{1}(u)$ again denotes the generating function of the excess degree distribution of the network. By inspection, $u=1$ is always a solution; this solution entails not having an infodemic. We are interested in whether or not there are also solutions $u\in(0,1)$.
In some special cases, it is possible to solve for $u$ analytically, but it is typically difficult (or even impossible) to do so even when $g_{1}(u)$ has a simple closed-form expression. However, it is effective to find the root $u^{*}\in(0,1)$ of section V.1 numerically using either an explicit expression or an approximation of the generating function of the excess degree distribution Newman (2018). One then substitutes $u^{*}$ into eq. 10 to obtain an approximation of $S$. The solution $u^{*}$ yields an approximation of $S$ when there is an infodemic.
In fig. 5, we show an example where we compare our approximation of the total number of content shares ($SN$) to numerical simulations in configuration-model networks of three different sizes. For each size (200, 2400, and 4800 nodes), we generate 1000 configuration-model networks with degree sequences from a Poisson distribution with mean $\lambda=5$. We plot histograms of the total number of content shares, and we observe good agreement between our analytical and numerical results.
Figure 5: Histograms of the total number of content shares across 1000 trials for 200-node, 2400-node, and 4800-node configuration-model networks. The dashed vertical line indicates the number of content shares (i.e., the dissemination-tree sizes) that we predict from our analysis. In each realization, we generate an $N$-node configuration-model network with a degree sequence that we obtain from a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.2$. In each realization, we draw a new degree sequence and a new set of node opinions.
VI Summary statistics for dissemination trees
Figure 6: The effect of varying the network size $N$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate an $N$-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. We vary $N$ from $100$ to $5000$ in increments of $100$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. Both the total number of content shares (i.e., the dissemination-tree size) and the width grow linearly with $N$. The longest-path length grows quickly at first as we increase $N$, and then it grows much more slowly with $N$. The structural virality appears to saturate at a constant value for sufficiently large $N$.
Figure 7: The effect of varying the expected mean degree $\lambda$ of a network on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda$. We vary $\lambda$ from $1$ to $40$ in increments of 1. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. The total number of content shares grows initially and then saturates at about $800$. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality increase at first and then decrease, eventually leveling off at a constant value for sufficiently large expected mean degrees. In our simulations, the longest-path length has a maximum at about $\lambda=4$ and structural virality has a maximum at about $\lambda=3$.
Figure 8: The effect of varying the receptiveness parameter $c$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. We vary $c$ from $0.02$ to $0.48$ in increments of $0.02$. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$. The total number of content shares (i.e., the dissemination-tree size) and the width are very small until about $c=0.1$, and then they grow with $c$. Both growth rates seem roughly linear, although the curve for the total number of content shares seems to be concave down and the receptiveness experiences some slower growth with $c$ before settling into linear growth. The longest-path length increases initially with $c$ before seeming to reach a maximum and then decaying. The structural virality also appears to increase initially with $c$ before seeming to reach a maximum and then decaying to a constant value. The standard deviations of the longest-path length and the structural virality are large.
Figure 9: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks with initial node opinions from a uniform distribution. We vary $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. The solid curves show the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. When $x_{0}\in(0.2,0.8)$, each of these four summary statistics is roughly constant. This is not surprising because we draw initial node opinions uniformly at random, so the transmission probability for content to spread along an edge depends only on $c$ in this interval (see eq. 4). The symmetry of these summary statistics with respect to $x_{0}$ is also clear from eq. 4. Content states that are closer to the boundary produce dissemination trees with fewer total content shares, smaller widths, and shorter longest paths. However, the structural virality of such “extreme” content states is slightly larger.
Figure 10: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks with initial node opinions that we draw from a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. We vary $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. The solid curves show the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$.
In this section, we perform numerical experiments and study the output of our bounded-confidence content-spread model on configuration-model networks. We quantify its features by computing four summary statistics of the dissemination trees that we obtain in our simulations.
  • Total number of content shares. One way to measure the effectiveness of content spread is to count the total number of nodes that spread (i.e., “adopt”) the content. This quantity, which we studied in section V, is equal to the number of nodes in the dissemination tree $G_{0}$.
  • Length of a longest adoption path. An adoption path is a path in a dissemination tree from the source node to another node in the tree Oh and Porter (2018). We measure the “depth” of content spread of the content by calculating the length of a longest adoption path.
  • Width. One can arrange a dissemination tree $G_{0}$ with the source node at the top (i.e., at level 0), nodes with adoption-path length 1 at level 1, nodes with adoption-path length 2 at level 2, and so on. The width of a dissemination tree is the largest number of nodes that adopt the content in a level. Mathematically, the width is
    $\max_{a\in\{1,\ldots,l_{*}\}}\text{(number of nodes with adoption path length }a)\,,$
    where $l_{*}$ is the length of the longest adoption path.
  • Structural virality. Structural virality (i.e., the Wiener index) was used by Goel et al. Goel et al. (2016) to measure the viral nature of content spread. The structural virality $v$ is the mean shortest-path length between nodes in the dissemination tree $G_{0}$. Mathematically, the structural virality is
    $v=\frac{1}{n(n-1)}\sum_{i=1}^{N}\sum_{j=1}^{N}d_{ij}\,,$
    where $d_{ij}$ is the length of a shortest path between nodes $i$ and $j$.
One natural question to ask is how the number of nodes of a network affects the spread of content on that network. We study this question by increasing the network size $N$ for fixed expected mean degree $\lambda$ (see fig. 6). With this construction, the expected number of edges remains constant as we increase $N$. As implied by our analysis in section V, the total number of content shares grows linearly with $N$ (with a slope indicated by eq. 10). The dissemination-tree width also appears to grow linearly with $N$. The longest-path length and structural virality change more dramatically for small $N$ than for large $N$. The former appears to grow very slowly for large $N$, and the latter appears to eventually saturate at a constant value.
In fig. 7, we examine the impact of increasing the expected mean degree $\lambda$ on the four summary statistics. The total number of content shares increases with $\lambda$ before saturating at a constant value. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality have maxima for small values of $\lambda$. When $\lambda$ is large, it is seemingly possible to for the content to be adopted by all receptive and connected nodes in just a few steps, yielding a large dissemination tree that reaches many new nodes in a single time step, short adoption paths, and small structural viralities. The maxima in the longest-path length and structural virality indicate that a spread process persists longer (and hence travels farther from the source node) when nodes have fewer neighbors on average (but still have large enough degree for an infodemic to occur).
Varying the receptiveness parameter $c$ (see fig. 8) yields similar dissemination-tree statistics as varying the expected mean degree. We again observe maxima in the longest-path length and the structural virality, likely due to larger receptiveness values favoring faster content spread, although these statistics have large standard deviations in this experiment. We observe evidence of a phase transition in the total number of content shares at the value of $c$ that we predicted on our analysis in section IV.1. There also appears to be an accompanying phase transition in the dissemination-tree width.
We also examine the impact of the content state on our dissemination-tree statistics for two different initial opinion distributions (see eq. 3). In fig. 9, we show results of simulations in we draw the initial node opinions from a uniform distribution on the interval $(0,1)$. In fig. 10, we show results of simulations in which we draw the initial node opinions on $\mathbb{R}$ from a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. As expected, we observe smaller values of all four summary statistics when the content state lies in the tails of a opinion distribution $\phi(x)$ (or near the boundary of the distribution interval for the uniform distribution). We observe maxima in the total number of content shares and dissemination-tree width at the distribution mean (and also for a very large interval around it for the uniform distribution). For the Gaussian distribution, the content states that yield the largest longest-path lengths and structural viralities occur away from the mean (and are symmetric), indicating that content can spread farther from the source node in these situations then when the content state equals the mean initial opinion. These dissemination trees are longer and narrow than when the content state equals the mean initial opinion. If the content state is too extreme, it spreads very little.
VII Conclusions and discussion
We examined “infodemics”, in the form of large content-spreading cascades, in a bounded-confidence content-spread model on configuration-model networks. To do this, we defined an “opinion reproduction number” that is analogous to the basic reproduction number from disease dynamics. By examining whether opinion spread is below or above the opinion reproduction number — such that it is below or above a corresponding “infodemic threshold” — we investigated when opinions die off and when they propagate widely in a population of agents. We quantified the size and structure of content spread by measuring several properties of dissemination trees — tree size, tree width, the length of the longest adoption paths, and structural virality — and we thereby illustrated how network structure and spreading-model parameters affect how content spreads. We found that larger networks, larger expected mean degrees, larger receptiveness, and content states near the mean of opinion distributions all promote larger total numbers of content shares, including the possibility of sharing the content to many new nodes in a single time step. Additionally, when the expected mean degree and receptiveness levels are small but place the spreading process above the infodemic threshold, there can be longer dissemination trees than when the expected mean degree or receptiveness are larger, even though the total number of content shares is smaller. This indicates that content that is spreading farther but less widely from a source agent (i.e., the agent that originally posts a piece of content).
There are many interesting ways to extend our work. A particularly relevant one is to examine the effects of purposeful choices of source nodes, such as to try to promote influence maximization Peng et al. (2018). One can also examine the effects of competing social contagions Gleeson et al. (2014), which perhaps are spread by agents with different political perspectives or agendas. One can also incorporate content mutation Adamic et al. (2016) (including through opinion averaging, as in classical BCMs such as the Deffuant–Weisbuch model Deffuant et al. (2000)) and examine how that changes infodemic thresholds and spreading patterns. Another worthwhile future direction is to adapt the recently developed “distributed reproduction numbers” She, Paré, and Hale (2023) from disease spreading to content spreading. It is also desirable to extend our study of content spreading to other network structures. One key direction is to consider more complicated network models (such as generalizations of configuration models that incorporate various types of heterogeneities Melnik et al. (2014)) and real-world networks. It is also natural to extend our investigation to more complicated types of networks, such as multilayer networks (which allow multiple types of social connections and communication channels) Kivelä et al. (2014), hypergraphs (which allow simultaneous interactions between three or more agents) Battiston et al. (2020), and adaptive networks (e.g., to incorporate changing relationship structures such as “unfollowing” or “unfriending” on social-media platforms) Berner et al. (2023).
Software
We performed computations and constructed visualizations using MATLAB. The code to reproduce our numerical simulations is available at https://github.com/hzinnbrooks/bounded-confidence-spreading-process.
Acknowledgements.
We thank Jonas Juul for helpful discussions. HZB was funded in part by the National Science Foundation (grant number DMS-2109239) through their program on Applied Mathematics. MAP acknowledges financial support from the National Science Foundation (grant number 1922952) through their program on Algorithms for Threat Detection (ATD). MAP thanks David Campbell for his friendship and many discussions over the years.
VIII A few words about David Campbell (by Mason A. Porter)
It is a great honor to contribute to Chaos’s special issue in celebration of David Campbell’s 80th birthday. David is one of my favorite people in science, and I deeply appreciate being asked to contribute an article to this.Festschrift. David Campbell’s research contributions and, especially, his foundational role in Chaos are well-known. I could write about that, but I will leave that to others. What is especially important is that David is a great human being. He has always been fair to me and he has been very good to me, as he has been to so many others.
I first interacted with David when I was a postdoc at Georgia Tech. I was naively trying to publish a modified version of my doctoral thesis Porter (2002) as a review article, and I submitted it to the journal Physics Reports. David, who I had yet to meet in person, was the handling editor of my submission. My paper was rejected, but what stood out in that process was the carefulness and decency in which David handled the process. All I want with a manuscript submission — and all anyone should ever want with a manuscript submission, as both a necessary and sufficient condition — is to be treated fairly and with dignity. David saw a teaching opportunity, and he could see that I had things to learn, and he went beyond the call of duty and even asked the referee to disclose his identity to help with that teaching. By contrast, as we all see repeatedly, many journal editors seem to be bean counters, rather than actual editors. It wasn’t my ideal outcome, but it was the correct outcome, and I learned a lot from it. In my own editing roles, I have tried to draw on the approach and lessons that David exemplified in this story.
David continued looking out for me, as he has looked out for many others. Having — I think? — still not yet met in person, David invited me to give a talk as part of a special session at the 2005 American Physical Society (APS) March Meeting about the celebrated Fermi–Pasta–Ulam–Tsingou (FPUT) problem and contribute an article about the dynamics of Bose–Einstein condensates Porter et al. (2005) to a special issue in Chaos about the FPUT problem. David’s talk had a pop-music reference (which was about the Eagles, if memory serves), and he mentioned that he figured that I’d appreciate the music reference. (He was right.) Four of us in the special session eventually coauthored a popular article about the FPUT problem Porter, Zabusky, and Campbell (2009). I eventually wrote a variety of papers on the FPUT problem (and on other related problems), including one Nelson, Porter, and Choubey (2018) that turned out to be refereed by David and one of his students. It was a particularly fair and thorough referee report — of the type that scientists dream of getting that indicates that sometimes the reviewing process actually works the way that it it supposed to — that my coauthors were spending time trying to address. I remember telling David by e-mail during our revision process to tell him about our paper because I figured that he would be interested in this paper, which concerns heterogeneities in FPUT lattices. My e-mail to David either was about something else entirely or was because I had just seen a paper of his on a related topic. In any event, without any idea that David was the referee, I told him in my e-mail about our paper and that we were currently dealing with a tough-but-fair referee report. I found out from David’s response that he and his student had written that report. Once again, David was doing things in a way that helps authors while upholding rigorous standards. In other words, David once again did things in the way that they are supposed to be done
A recurring theme in my stories is David looking out for a junior scientist, fair treatment while retaining rigor, and doing things the right way in scientific editing and review. We all have reviewing horror stories and publishing can cause many frustrations. But some scientists consistently do things the right way, and David is one of them.
The best complement that somebody like me who comes from Jewish heritage can pay to another person is to call them a “mensch” (which, essentially, is calling somebody a decent human being, as they are a person of integrity and honor). David Campbell is a mensch.
Happy birthday, David!